OpenStudy (calculusxy):

How do you figure out the percentage of a number? Please tell me the procedure.

OpenStudy (goldphenoix):

Give me an example, please.

OpenStudy (anonymous):

question- what is 25% of 100? answer- take the percentage and move the decimal 2 places to the left. so 25% would become (.25). then multiply that by the number you are trying to find the percentage of (100) so .25*100= 25

OpenStudy (calculusxy):

@GoldPhenoix Ok. So what about if I have to find 20% of £40.

OpenStudy (anonymous):

so it would be .20*40=8%

OpenStudy (anonymous):

sorry, exclude the percentage sign at the end of the 8

OpenStudy (goldphenoix):

Alright. Then you would want to use a proportion to solve this \[\large \large \frac{ is }{ of } = \frac{ percent }{ 100 }\] Now plug it in: \[\large \large \frac{ is }{ 40 } = \frac{ 20 }{ 100 }\] Now cross multiply. :) 40 * 20 = 800 800 = 100x Divide both by 100 to get 8.

OpenStudy (goldphenoix):

I hope this help! :)

OpenStudy (calculusxy):

Sorry for making this complicated, but if I do 40% of £20 then it is still the same. I don't understand that. Some people are saying that it is £32.

OpenStudy (calculusxy):

Meaning that it is still £8.

OpenStudy (goldphenoix):

What are you trying to say? o.o.

OpenStudy (calculusxy):

I got confused from this. Please can you clarify this? "If a product was 20% off of $40 it would cost $32. ($40 minus $8 equals $32)" from the Ask.com community.

OpenStudy (goldphenoix):

Correct.

OpenStudy (calculusxy):

Can you explain it to me?

OpenStudy (goldphenoix):

Explain to you about what?

OpenStudy (calculusxy):

About the figuring of 32 from the citing posted above.

OpenStudy (goldphenoix):

Oh. So there's a product that cost 40 dollars. And that product was on sale for 20%. So you want to multiply 80% or .8 by 40 because you are still paying 80% of the original price. ;) So 40 * .8 = 32. I hope this help you understand! :)

OpenStudy (calculusxy):

Thank you.